# Adjoint log canonical foliated singularities on surfaces

### v1

@@ Abstract.
Let $(X,\mathcal{F})$ be a foliated surface over $\mathbb{C}$. We study the singularities of the adjoint foliated divisor $K_{\mathcal{F}}+\epsilon K_{X}$. We provide a complete classification of $\epsilon$-adjoint log canonical singularities of foliated surfaces for $\epsilon\in(0,1/3)$. Moreover, we prove that for any $\epsilon\in(0,1/5)$, every $\epsilon$-adjoint log canonical singularity is log canonical for $\mathcal{F}$, and that for any $\epsilon\in(0,1/4)$, every $\epsilon$-adjoint canonical singularity is log canonical for $\mathcal{F}$. Finally, we present examples showing that both bounds are sharp.
@@ 1. Introduction
The study of singularities lies at the core of algebraic geometry. Singular varieties arise naturally, for instance, in the minimal model program and in the construction of moduli spaces. Among the most fundamental classes are (log) canonical singularities, which have been extensively studied, particularly in dimension two. In fact, log canonical surface singularities admit a complete classification (see [Alex92], [KM98, Sec. 4.1], or Theorem A.8).
In recent years, the theory of foliations in algebraic geometry has attracted increasing attention. It is therefore natural to seek foliated analogues of the classical theory of singularities. In this direction, McQuillan [McQ08] introduced the notion of (log) canonical singularities for foliations (cf. Definition 2.12), and obtained a classification of canonical foliated singularities on surfaces (see [McQ08, Corollary I.2.2, Fact I.2.4, Theorem III.3.2]). Subsequently, Y. Chen classified log canonical foliated singularities on surfaces (see [Chen23, Theorem 1.1] or Theorem B.9).
More recently, Pereira and Svaldi introduced adjoint foliated structures in [PS19]. This framework was further developed by Spicer and Svaldi for rank-one foliations on surfaces [SS23], and by Cascini–Han–Liu–Meng–Spicer–Svaldi–Xie for algebraically integrable foliations on higher-dimensional varieties [CHL+24]. Adjoint foliated structures have proved effective in addressing several difficulties in the birational geometry of foliations. For instance, effective birationality fails even for rank-one foliations on surfaces [Lu25]. The underlying idea is that, instead of studying the canonical divisor $K_{\mathcal{F}}$ alone, one considers divisors of the form $K_{\mathcal{F}}+\epsilon K_{X}$ for $0<\epsilon\ll 1$, which allows one to exploit classical results for $K_{X}$.
In this paper, we study singularities of adjoint foliated structures $K_{\mathcal{F}}+\epsilon K_{X}$ on surfaces. Our goal is to understand how $\epsilon$-adjoint singularities relate to the usual singularities of foliations. More precisely, we consider the following questions.
  • Q1.
    Does there exist $\epsilon_{0}>0$ such that $\epsilon$-adjoint canonical singularities are canonical for all $\epsilon\in(0,\epsilon_{0})$?
  • Q2.
    Does there exist $\epsilon_{0}>0$ such that $\epsilon$-adjoint log canonical singularities are log canonical for all $\epsilon\in(0,\epsilon_{0})$?
We first observe that Question (Q1) admits a negative answer (see Example 2.15). This shows that $\epsilon$-adjoint canonical singularities need not be canonical even for $\epsilon>0$ sufficiently small.
Motivated by this failure of Question (Q1), we are led to consider the following weaker variant.
  • Q3.
    Does there exist $\epsilon_{0}>0$ such that $\epsilon$-adjoint canonical singularities are log canonical for all $\epsilon\in(0,\epsilon_{0})$?
The main purpose of this paper is to address these questions. We give a complete classification of $\epsilon$-adjoint log canonical singularities of foliations for $\epsilon\in(0,1/3)$, and prove that Questions (Q2) and (Q3) admit affirmative answers.
We now state our main results.
Theorem 1.1 (cf. Theorem 4.14).
@@
Fix $\epsilon\in(0,1/5)$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface, and assume that $p$ is an $\epsilon$-adjoint log canonical singularity of $(Y,\mathcal{G})$. Let
@@
be the minimal resolution (cf. Definition 2.5) of $(Y,\mathcal{G},p)$, with exceptional divisor $E=\bigcup_{i=1}^{r}E_{i}$. Then $E$ belongs to one of the following families:
@@
    Chains of $\mathcal{F}$-invariant curves: either an $\mathcal{F}$-chain (Figure 3), two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail (Figure 4), or a chain of (2)(-2)-$\mathcal{F}$-curves (Definition 2.18).
@@
    $\mathcal{F}$-dihedral graphs: two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail, which in turn is attached to a chain of (2)(-2)-$\mathcal{F}$-curves (see Figure 5).
@@
    Elliptic Gorenstein leaves: either a rational $\mathcal{F}$-invariant curve with a single node, or a cycle of (2)(-2)-$\mathcal{F}$-curves (see Figure 6).
  4. (4)
    Smooth elliptic non-$\mathcal{F}$-invariant curves: a smooth elliptic non-$\mathcal{F}$-invariant curve $E_{1}$ with ${\rm tang}(\mathcal{F},E_{1})=0$.
  5. (5)
    $\mathcal{F}$-star chains: a chain containing a unique smooth rational non-$\mathcal{F}$-invariant curve $E_{k}$ with ${\rm tang}(\mathcal{F},E_{k})=0$, while the remaining components form $\mathcal{F}$-chains (see Figure 8).
  6. (6)
    $\mathcal{F}$-star graphs: a smooth rational non-$\mathcal{F}$-invariant curve $C$ with ${\rm tang}(\mathcal{F},C)=0$ at the center, together with attached $\mathcal{F}$-chains of type $(l;m_{1},\dots,m_{l})$, where $(l;m_{1},\dots,m_{l})$ is one of the following:
    1. (a)
      $l=3$, $(m_{1},m_{2},m_{3})=(2,2,n)$ for $n\geq 2$, or $(2,3,3)$, $(2,3,4)$, $(2,3,5)$;
    2. (b)
      $l=3$, $(m_{1},m_{2},m_{3})=(2,3,6)$, $(2,4,4)$, or $(3,3,3)$;
    3. (c)
      $l=4$, $(m_{1},\dots,m_{4})=(2,2,2,2)$.
    (See Definition 2.22 and Figure 7.)
Moreover, configurations of type (3), (4), (6-b), and (6-c) are not $\epsilon$-adjoint klt.
@@
For $\epsilon\in[1/5,1/3)$, there are additional configurations: (cf. Theorem 4.14)
  1. (7)
    A rational smooth $\mathcal{F}$-invariant curve $C$ with ${\rm Z}(\mathcal{F},C)=3$ connecting to two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$ of type $(2,3)$.
  2. (8)
    A rational smooth $\mathcal{F}$-invariant curve $C$ with ${\rm Z}(\mathcal{F},C)=3$ connecting to two $\mathcal{F}$-chains $\Theta_{1}$ and $\Theta_{2}$ of type $(2,3)$ and a chain of (2)(-2)-$\mathcal{F}$-curves.
Here, by an $\mathcal{F}$-chain $\Theta$ of type $m$ we mean that $m$ denotes the absolute value of the determinant of the intersection matrix of $\Theta$.
Remark 1.3.
For $\epsilon\in(0,1/4)$, cases (3), (4), (6), (7) and (8) fail to be $\epsilon$-adjoint canonical (see Theorem 5.6).
Therefore, we have the following proposition.
@@
Let $(X,\mathcal{F})$ be a foliated surface.
  1. (1)
    For any $\epsilon\in(0,1/5)$, if $p$ is an $\epsilon$-adjoint log canonical singularity of $(X,\mathcal{F})$, then $p$ is a log canonical singularity of $\mathcal{F}$. Moreover, $p$ is a log canonical singularity of $X$. If furthermore $p$ is $\epsilon$-adjoint klt, then $p$ is a klt singularity of $X$.
  2. (2)
    For any $\epsilon\in(0,1/4)$, if $p$ is an $\epsilon$-adjoint canonical singularity of $(X,\mathcal{F})$, then $p$ is a log canonical singularity of $\mathcal{F}$ and a klt singularity of $X$.
Proof.
This is a direct consequence of Theorem 4.14, Theorem 5.6 and Theorem A.8. ∎
Remark 1.5.
  1. (1)
    Proposition 1.4(1) answers a question posed to the author by Professor Jihao Liu. In [SS23], the corresponding result was proved under the assumption that $p$ is smooth. In [Vas25], it was proved under the assumption that $p$ is a klt singularity.
  2. (2)
    The bounds in Proposition 1.4 are both sharp, see Example 2.16.
As a direct corollary of Proposition 1.4(2), we obtain the following result.
Corollary 1.6.
Let $X$ be a smooth projective surface and $\mathcal{F}$ a rank one foliation with canonical singularities. Then for any $\epsilon\in(0,1/4)$, there exists a birational morphism $\varphi:X\to Y$ (which is a $(K_{\mathcal{F}}+\epsilon K_{X})$-non-positive contraction) such that either
@@
    there exists a morphism $f:Y\to Z$ such that $\rho(Y/Z)=1$ and $(K_{\mathcal{G}}+\epsilon K_{Y})$ is $f$-ample.
Moreover, $Y$ has klt singularities and $\mathcal{G}$ has log canonical singularities.
Proof.
This directly follows from Proposition 1.4(2) and [KM98, Corollary 3.17]. ∎
@@
Corollary 1.6 improves [SS23, Theorem 1.1], where the authors treat the case $\epsilon\in(0,1/5)$. It also provides a useful complement to [LWX25, Theorem 1.1].
@@ 2.1. Foliations and singularities
    The singularity $p$ is called reduced if one of the two eigenvalues, say, $\lambda_{2}$, is nonzero and and the quotient $\lambda=\lambda_{1}/\lambda_{2}$ is not a positive rational number. In particular, if $\lambda=0$, we call $p$ a saddle-node.
@@ 2.2. Minimal resolution of foliated surfaces
Let $(X,\mathcal{F})$ be a foliated surface. Let $p$ is a singularity of $(X,\mathcal{F})$. A resolution of $(X,\mathcal{F},p)$ is a birational morphism
@@
    for any $q\in\pi^{-1}(p)$, $p$ is a smooth point of $Y$ and $p$ is a reduced (or regular) point of the induced foliation $\mathcal{G}:=\pi^{*}\mathcal{F}$.
@@ 2.3. Index Theorems
A curve $C\subseteq X$ is said to be $\mathcal{F}$-invariant if the inclusion $T_{\mathcal{F}}|_{C}\rightarrow T_{X}|_{C}$ factors through $T_{C}$, where $T_{C}$ is the tangent bundle of $C$.
@@ 2.3.2. Invariant curves
By definition, ${\rm Z}(\mathcal{F},C,p)={\rm CS}(\mathcal{F},C,p)=0$ if $p$ is not a singularity of $\mathcal{F}$. If $p$ is a reduced singularity of $\mathcal{F}$, then ${\rm Z}(\mathcal{F},C,p)\geq 0$. Let
@@
    ${\rm CS}(\mathcal{F},x=0,p)=a,\quad\text{and}\quad{\rm Z}(\mathcal{F},x=0,p)=k+1.$
@@ 2.4. Pairs and Triples
Let $X$ be a normal surface and let $\mathcal{F}$ be a foliation on $X$. Let $D$ be an $\mathbb{R}$-divisor on $X$. Then we can uniquely decompose $D$ as

### v2

# Sharp Thresholds for $\epsilon$-Adjoint Singularities of Foliated Surfaces

@@ Abstract.
Let $(X,\mathcal{F})$ be a foliated surface over the complex numbers. We study the variation of $\epsilon$-adjoint singularities, defined by the adjoint divisor $K_{\mathcal{F}}+\epsilon K_{X}$ ($\epsilon>0$), and analyze their stability as $\epsilon$ varies.
We prove that a sharp first stability threshold occurs at $\epsilon=1/5$: for $\epsilon\in(0,1/5)$, every $\epsilon$-adjoint log canonical singularity is foliated log canonical, while at $\epsilon=1/5$ a boundary configuration enters the admissible region. In the adjoint canonical setting, the maximal stability interval is $\epsilon\in(0,1/4)$. Both thresholds are optimal and arise from explicit extremal configurations.
These results are obtained via a complete classification of $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$ in terms of negative definite exceptional configurations.
@@ 1. Introduction
Singularities play a central role in birational geometry, arising naturally in the minimal model program and in the construction of moduli spaces. Among them, (log) canonical singularities form one of the most fundamental classes. In dimension two, log canonical surface singularities admit a complete classification (see [Alex92], [KM98, Sec. 4.1], or Theorem 2.24).
For foliations, McQuillan [McQ08] introduced (log) canonical singularities of foliations and classified canonical singularities of foliations on surfaces. The log canonical case was later established by Y. Chen [Chen23, Theorem 1.1] (see also Theorem 2.25). These results provide a detailed structural description of log canonical singularities of foliations on surfaces.
Adjoint foliated structures, introduced by Pereira and Svaldi [PS19], interpolate between the intrinsic birational geometry of a foliation and the ambient geometry of the surface. This framework was further developed by Spicer and Svaldi for rank-one foliations on surfaces [SS23], and by Cascini–Han–Liu–Meng–Spicer–Svaldi–Xie for algebraically integrable foliations in higher dimensions [CHL+24]. Adjoint divisors
$K_{\mathcal{F}}+\epsilon K_{X},\qquad 0<\epsilon\ll 1,$
provide a natural deformation of the foliated canonical class and address structural pathologies in the birational theory of foliations. For instance, the canonical ring of a foliation need not be finitely generated even if $K_{\mathcal{F}}$ is big and nef (cf. [McQ08, Theorem IV.2.2]), and effective birationality may fail even for rank-one foliations [Lu25].
A natural question is whether sufficiently small adjoint perturbations preserve canonical singularities. However, this is not always the case: there exist adjoint canonical singularities for which arbitrarily small adjoint deformations may destroy canonicity (see Example 6.1). Thus the canonical condition is not stable under small variations of $\epsilon$.
Since canonical singularities form a subclass of log canonical ones, we first study adjoint log canonical singularities, with adjoint canonical singularities imposing additional constraints.
Passing to a minimal resolution, the $\epsilon$-adjoint log canonical singularities correspond to rational polyhedral regions defined by a finite system of linear inequalities, whose coefficients depend on $\epsilon$ and on the configuration of exceptional curves. In this way, the variation of $\epsilon$ can be interpreted as a linear deformation of discrepancy inequalities. As $\epsilon$ increases, the polyhedral region shrinks linearly, and critical values
$0<\epsilon_{1}<\epsilon_{2}<\cdots$
occur when one inequality becomes sharp, giving rise to combinatorial wall-crossing phenomena at which new configurations may enter the admissible region. Our primary focus is on the first critical value $\epsilon_{1}$, which governs the initial stability behavior of adjoint singularities.
In dimension two, the geometry of negative definite configurations allows these extremal situations to be tracked explicitly. To study $\epsilon_{1}$, it is convenient to work with a slightly larger interval
$\epsilon\in(0,1/3),$
where the upper bound $1/3$ is dictated by the structure of the discrepancy inequalities, marking the next potential threshold beyond which new configurations may appear. Within this interval, the $\epsilon$-adjoint log canonical singularities are controlled by a finite set of structural types, which we classify completely below.
Theorem 1.1 (Main classification theorem).
@@
Fix $\epsilon\in(0,1/3)$. Let $(Y,\mathcal{G},p)$ be a germ of a foliated surface such that $p$ is $\epsilon$-adjoint log canonical. Let
@@
be the minimal resolution of $(Y,\mathcal{G},p)$, and write $E=\bigcup_{i=1}^{r}E_{i}$ for the exceptional divisor.
Then $E$ belongs to one of the following structural types:
@@
    Foliated canonical configurations. All components are $\mathcal{F}$-invariant, including $\mathcal{F}$-chains, two (1)(-1)-$\mathcal{F}$-curves of self-intersection 2-2 joined by a bad tail, chains of (2)(-2)-$\mathcal{F}$-curves, $\mathcal{F}$-dihedral forks, and elliptic Gorenstein leaves.
@@
    Foliated lc (non-canonical) configurations. There exists a smooth non-$\mathcal{F}$-invariant curve $E_{0}$ with ${\rm tang}(\mathcal{F},E_{0})=0$, and all other components (if any) form $\mathcal{F}$-chains attached to $E_{0}$. The curve $E_{0}$ is either rational (with attached chains) or elliptic (in which case $E=E_{0}$).
@@
    Boundary configurations. A smooth rational $\mathcal{F}$-invariant curve $C$ with ${\rm Z}(\mathcal{F},C)=3$ connecting two $\mathcal{F}$-chains of type $(2,3)$, possibly together with an additional chain of (2)(-2)-$\mathcal{F}$-curves. (Here, by an $\mathcal{F}$-chain $\Theta$ of type $m$ we mean that $m$ denotes the absolute value of the determinant of the intersection matrix of $\Theta$.)
Configurations of type (1) and (2) are precisely the foliated log canonical singularities. Type (3) is not foliated log canonical. It becomes $\epsilon$-adjoint log canonical precisely at $\epsilon=1/5$, and does not occur for $\epsilon<1/5$.
A complete classification, with all subtypes and details, is given in Theorem 4.15.
1.1. Log canonical stability
The boundary configuration enters the admissible region precisely at $\epsilon=1/5$, marking the first wall in the $\epsilon$-direction and the first stability transition in the adjoint log canonical setting. Below this value, adjoint log canonicity forces foliated log canonicity, whereas at $\epsilon=1/5$ the boundary configuration becomes admissible.
Proposition 1.2.
[Corollary 4.16+Example 6.3] Fix $\epsilon\in(0,1/5)$. Suppose $p$ is an $\epsilon$-adjoint log canonical singularity of $(X,\mathcal{F})$. Then:
  • $p$ is log canonical both as a foliated singularity and as a surface singularity;
  • moreover, if $p$ is $\epsilon$-adjoint klt, then $p$ is klt as a singularity of $X$.
The threshold $\epsilon=1/5$ is sharp: at this value, the $\epsilon$-adjoint log canonical foliated singularity may fail to be log canonical as a foliated singularity.
@@ 1.1. Log canonical stability
Proposition 1.2 extends previous results: [SS23, Lemma 2.19] treats the case where $p$ is smooth, and [Vas25, Lemma 2.13] considers the case where $p$ is klt as a surface singularity.
1.2. Log-to-canonical stability
In contrast to the log canonical case, imposing the adjoint canonical condition further restricts the admissible configurations. The boundary configuration does not occur if $\epsilon<1/4$ and the singularity is $\epsilon$-adjoint canonical. Thus $\epsilon=1/4$ marks the first wall in the adjoint canonical setting. Below this value, adjoint canonicity forces foliated log canonicity, and the admissible configurations form a strictly smaller class, which we classify completely (see Theorem 5.7).
Analogous to Proposition 1.2, we also have the following log-to-canonical stability proposition.
@@ 1.2. Log-to-canonical stability
[Corollary 5.9+Example 6.3] Fix $\epsilon\in(0,1/4)$. Suppose $p$ is an $\epsilon$-adjoint canonical singularity of $(X,\mathcal{F})$. Then $p$ is a log canonical singularity of $\mathcal{F}$ and a klt singularity of $X$.
The threshold $\epsilon=1/4$ is sharp: at this value, the $\epsilon$-adjoint canonical foliated singularity may fail to be log canonical as a foliated singularity.
1.3. Applications to the adjoint minimal model program
As a direct consequence of Proposition 1.4, we obtain an adjoint minimal model program in the range $\epsilon\in(0,1/4)$. In particular, the threshold $1/4$ determines precisely the maximal range in which the adjoint MMP preserves the expected singularity behavior.
Proposition 1.5.
[See Section 5.1] Let $X$ be a smooth projective surface and $\mathcal{F}$ a rank one foliation with canonical singularities. Then for any $\epsilon\in(0,1/4)\cap\mathbb{Q}$, there exists a birational morphism $\varphi:X\to Y$ (which is a $(K_{\mathcal{F}}+\epsilon K_{X})$-non-positive contraction) such that either
@@ 1.3. Applications to the adjoint minimal model program
    there exists a morphism $f:Y\to Z$ such that $\rho(Y/Z)=1$ and $(K_{\mathcal{G}}+\epsilon K_{Y})$ is $f$-ample,
where $Y$ has klt singularities and $\mathcal{G}$ has log canonical singularities. Moreover, if $K_{\mathcal{F}}+\epsilon K_{X}$ is big, then $K_{\mathcal{G}}+\epsilon K_{Y}$ is ample.
@@
Proposition 1.5 extends [SS23, Theorems 1.1 and 1.2], where the case $\epsilon\in(0,1/5)$ was treated. It also complements [LWX25, Theorem 1.1].
1.4. Interpolated lc thresholds and gap phenomena
The threshold $1/5$ admits a natural interpretation in terms of the interpolated lc threshold, introduced by McKernan [McK22] (see also [CHL+24, Appendix A]).
For a foliated surface $(X,\mathcal{F})$, define
$t_{0}:=\sup\left\{\,t\in[0,1]\mid tK_{\mathcal{F}}+(1-t)K_{X}\text{ is lc}\,\right\}.$
Noting that $K_{\mathcal{F}}+\epsilon K_{X}$ corresponds to the above expression via $\epsilon=\frac{1-t}{t}$ for $t>0$, Proposition 1.2 immediately implies:
Proposition 1.7.
The interpolated lc threshold $t_{0}$ satisfies
$t_{0}=1\quad\text{or}\quad t_{0}\leq\frac{5}{6}.$
More generally, the following 1-gap conjecture for interpolated lc thresholds was formulated in [CHL+24, Conjecture A.6]:
Conjecture 1.8 (1-gap for interpolated lc thresholds).
Let $d$ be a positive integer. Then there exists a positive real number $\tau=\tau(d)$ satisfying the following property:
Let $(X,\mathcal{F},t)$ be an adjoint foliated structure of dimension $d$. Assume that $tK_{\mathcal{F}}+(1-t)K_{X}$ is log canonical for some $t>1-\tau$ and that $K_{\mathcal{F}}$ is $\mathbb{Q}$-Cartier. Then $\mathcal{F}$ is log canonical.
In dimension 2, our results show that the constant is
$\tau(2)=\frac{1}{6},$
and that this value is sharp.
Outlook. Our results reveal a sharp first threshold for adjoint singularities in dimension two. It is natural to ask whether analogous threshold behavior occurs in higher dimensions, and how such gap principles may interact with boundedness, moduli problems, or the birational geometry of foliated pairs.
Organization of the paper. Section 2 reviews background on foliations and singularities. Section 3 establishes structural properties of negative definite configurations. Section 4 classifies $\epsilon$-adjoint log canonical singularities for $\epsilon\in(0,1/3)$ and determines the threshold $1/5$. Section 5 studies $\epsilon$-adjoint canonical singularities, determines the threshold $1/4$, and proves the adjoint minimal model program (MMP). Section 6 presents examples illustrating the sharpness of both thresholds.
@@ 2.1. Foliations and singularities
    The singularity $p$ is called reduced if one of the two eigenvalues, say $\lambda_{2}$, is nonzero and the quotient $\lambda=\lambda_{1}/\lambda_{2}$ is not a positive rational number. In particular, if $\lambda=0$, we call $p$ a saddle-node.
@@ 2.2. Minimal resolution of foliated surfaces
Let $(X,\mathcal{F})$ be a foliated surface and let $p$ be a singularity of $(X,\mathcal{F})$. A resolution of $(X,\mathcal{F},p)$ is a birational morphism
@@
    for any $\in\pi^{-1}(p), qq$ is a smooth point of $Y$ and $q$ is a reduced (or regular) point of the induced foliation $\mathcal{G}:=\pi^{*}\mathcal{F}$.
@@ 2.3. Index Theorems
A reduced curve $C\subseteq X$ is said to be $\mathcal{F}$-invariant if the inclusion $T_{\mathcal{F}}|_{C}\rightarrow T_{X}|_{C}$ factors through $T_{C}$, where $T_{C}$ is the tangent bundle of $C$.
@@ 2.3.2. Invariant curves
By definition, both ${\rm Z}(\mathcal{F},C,p)$ and ${\rm CS}(\mathcal{F},C,p)$ vanish if $p$ is not a singularity of $\mathcal{F}$. If $p$ is a reduced singularity of $\mathcal{F}$, then ${\rm Z}(\mathcal{F},C,p)\geq 0$. Let
@@
    ${\rm CS}(\mathcal{F},x=0,p)=a,\quad\text{and}\quad{\rm Z}(\mathcal{F},x=0,p)=k+1\,(\geq 2).$
@@ 2.4. Pairs and Triples
Let $X$ be a normal surface and let $\mathcal{F}$ be a foliation on $X$. Let $D$ be an $\mathbb{R}$-divisor on $X$. Then $D$ admits a unique decomposition